# Matching Non-Identical Objects

### v1

@@ 5.2 Robust Image Matching
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
Gaussian Noise 43.09 53.27 45.67 52.24 34.97
Shot Noise 43.41 53.82 46.10 51.72 40.00
Impulse Noise 44.98 50.11 44.91 50.05 37.73
Defocus Blur 32.69 48.35 45.58 49.23 57.22
Frosted Glass Blur 33.37 47.88 47.40 49.94 54.48
Motion Blur 42.12 53.63 44.76 51.84 55.60
Zoom Blur 24.40 31.21 22.87 30.97 23.08
Snow 31.51 30.33 25.26 23.21 31.85
Frost 32.24 31.71 26.39 24.22 19.23
Fog 70.99 73.49 71.19 73.89 67.10
Brightness 75.48 75.08 73.03 70.84 75.45
Contrast 39.50 38.47 39.47 42.61 53.22
Elastic Transform 54.78 66.21 60.37 64.36 58.56
Pixelate 67.94 68.24 65.01 64.81 74.77
Jpeg Compression 29.54 36.97 39.66 44.70 59.05
Average 44.40 50.58 46.51 49.64 49.49
Time per pair [msec.] 46.67 68.43 105.72 128.18 312.54
@@
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
None (Clean) 80.61 78.42 78.04 75.08 80.93
Gaussian Noise 27.54 41.01 41.93 44.98 33.02
Shot Noise 32.10 42.35 41.75 43.82 36.53
Impulse Noise 35.95 43.67 40.97 42.91 34.92
Defocus Blur 18.25 32.13 23.16 30.10 49.40
Frosted Glass Blur 34.25 44.80 42.96 46.21 54.53
Motion Blur 50.07 53.10 44.27 52.69 52.81
Zoom Blur 34.67 40.04 31.18 37.11 31.09
Snow 56.24 58.15 47.22 48.53 49.15
Frost 62.20 64.87 54.34 57.26 43.17
Fog 76.96 78.40 75.18 74.87 74.06
Brightness 77.14 76.98 76.78 75.06 78.41
Contrast 43.09 45.17 40.44 43.53 39.66
Elastic Transform 64.93 68.34 68.02 69.92 68.95
Pixelate 66.80 68.43 67.16 65.46 76.21
Jpeg Compression 47.41 53.82 60.39 60.12 68.14
Average 48.51 54.08 50.38 52.84 52.67
We next quantitatively evaluate the proposed method on a standard image matching task but with common corruptions. To our knowledge, we present the first evaluation of the robustness of image matching under common corruptions. We address two cases: (i) both input images are corrupted and (ii) only one of them is corrupted. The former shows the robustness against corruptions, and the latter shows the robustness against domain shift (i.e., clean image domain v.s. corrupted image domain).
Dataset.
We use image pairs from the MegaDepth-1500 [36] test set as the basis for various experiments. MegaDepth-1500 is a set of 1,500 image pairs from the MegaDepth [15] dataset for two popular phototourism destinations. MegaDepth consists of 3D reconstructed data from one million images, its camera poses and depth maps are first computed from COLMAP [31] and then refined as ground truth. Common Corruptions [13] simulates the 15 types of corruptions that occur when images are captured. The severity level of corruption is set to 5, which is the most challenging condition.
Metric.
[…]
Robustness against common corruptions.
Table 1 shows the AUC of matching when input images are both corrupted. The results of the dense matching method (i.e., LoFTR) is given for reference. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight increase in the runtime. The average AUC even exceeds that of LoFTR. The matching methods benefit from the proposed method particularly when images are corrupted by noises and blurs. We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method. These corruptions introduce edgy geometric patterns, and the matching methods find correspondence between them. See Section B.1 in the supplementary materials for more detailed comparisons.
Robustness against environmental changes.
As the objects to be matched are non-identical, they can appear in largely different environments. Table 2 shows the AUC when only either of the input images is corrupted and the other is clean. As in Tab. 1, the proposed method boosts the LightGlue and GlueStick for most cases and the average case. Interestingly, the results are better for almost all weather categories. For brightness, it is the only one of all categories that shows a decrease in accuracy, but the drop is slight. Weather categories such as Snow and Frost add strong geometry to the overall image, causing mismatches as occlusions. Heatmap weighting is considered to structurally complement the areas hidden by occlusion. The proposed method makes existing matchers robust to environmental changes in general, with a slight precision reduction in clean images. See Section B.2 in the supplementary material for more detailed comparisons.
5.3 Limitations
We here present failure cases. Figure 6 shows typical cases caused by (a) heatmaps not covering a whole part of objects and (b) great structural differences. In the former case, the importance scores take large only on particular parts of the objects. Further, the large importance scores on the irrelevant object disrupt the matching. This may be resolved by using a better visual explanation model [47, 32, 2, 1, 20, 34]. For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched. To address these cases, one may introduce recent vision and language models to complement the visual information with text features.
6 Conclusion
In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical as in prior studies. Because of the non-identicality, two objects may be captured in different environments and domains. Thus, this task also encompasses robust matching under image corruptions and domain shifts. The proposed method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects. Our experiments demonstrated the effectiveness of the proposed method in non-identical object matching and relative pose estimation under various image corruptions.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP22H03658 and JP22K17962. We thank Satomi Tanaka (Ricoh Company, Ltd.) for her fruitful comments and financial support from Ricoh Company, Ltd..
Supplementary Material
A More matching non-identical objects
In Sec. 5.1 of the main text, we evaluated non-identical object matching by inserting the proposed method between SuperPoint [7] and LightGlue [17]. The matching results for various image pair categories show that the proposed method enhances existing matchers and provides dense and consistent correspondences. To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$).
Setup.
We collected various images with objects from the COCO [16] and ImageNet [6] datasets. To prepare drawing or sketch versions of several images, we used the DreamStudio [^4] https://beta.dreamstudio.ai/generate, which is backended by the Stable Diffusion [28]. For details of the experiment, refer to Sec. 5.1.
A.1 Success cases
Figure 7 shows the matching results for the same class case. As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one. The combination of SuperPoint and LightGlue with our method is robust to object color and shape differences and performs dense and consistent matching. Even in the case of class discrepancy, as shown in Fig. 8, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals. Figure 9 shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings). These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
A.2 Failure cases
There are some cases of failure in non-identical object matching, as shown in Fig. 10. It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller), or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object. Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
Figure 7: Matching results in the same class case. Image pairs show objects of the same class in ImageNet [6] or COCO [16] datasets. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] only finds a small number of correspondence, and many of them are incorrect. (b) The proposed method significantly improves the matching, and performs dense and consistent correspondences.
Figure 8: Matching results in the class discrepancy case. Image pairs show objects with class discrepancies in ImageNet [6] or COCO [16] datasets. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] finds many inconsistent and scattered matches, and it is unclear which is the correct correspondence. (b) The proposed method finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
Figure 9: Matching results in the domain shift case. Image pairs for matching are ImageNet [6] or COCO [16] datasets with common corruptions [13], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [28]. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] finds correct matching, but also finds mismatches with many non-correspondence areas. (b) The proposed method reduces these mismatches and further improves to denser matching between objects.
Figure 10: Failure cases of non-identical object matching. (a) A combination of SuperPoint (SP) [7] and LightGlue (LG) [17] fails in matching. (b) The proposed method improves several matchings but still leaves many corresponding areas unmatched.
B Robust image matching details
[…]
B.1 Robustness against common corruptions.
[…]
B.2 Robustness against environmental changes.
[…]
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
None (Clean) 67.89 66.98 64.11 61.94 68.88
Gaussian Noise 27.44 37.28 28.64 35.11 20.07
Shot Noise 27.07 39.79 27.81 33.42 24.27
Impulse Noise 28.70 37.25 29.10 34.44 22.27
Defocus Blur 18.58 24.51 30.12 33.78 43.27
Frosted Glass Blur 17.24 27.61 29.16 31.75 39.21
Motion Blur 27.08 34.85 27.98 35.46 40.68
Zoom Blur 12.50 20.23 12.64 20.72 12.01
Snow 18.35 12.56 12.46 10.49 18.94
Frost 20.86 18.75 17.64 15.48 11.15
Fog 56.32 62.62 55.14 57.24 53.05
Brightness 61.10 63.95 58.75 56.51 61.88
Contrast 24.33 23.10 24.31 27.45 38.57
Elastic Transform 36.76 47.02 41.53 45.52 41.27
Pixelate 51.60 60.66 47.60 47.32 62.55
Jpeg Compression 16.60 27.73 22.35 27.64 43.49
Average 29.64 35.86 31.02 34.16 35.51
[…]
Gaussian Noise 14.70 24.44 14.63 21.53 9.85
Shot Noise 13.71 23.44 13.44 18.06 12.21
Impulse Noise 15.65 21.25 16.90 20.64 10.92
Defocus Blur 8.22 11.69 16.59 20.64 27.77
Frosted Glass Blur 7.17 18.09 14.58 16.12 24.59
Motion Blur 14.72 20.72 15.26 22.34 25.37
Zoom Blur 5.38 8.82 6.02 14.16 5.17
Snow 9.48 8.32 5.92 3.87 9.91
Frost 11.92 11.29 9.92 7.72 6.09
Fog 39.01 43.43 35.88 37.58 36.94
Brightness 43.60 47.98 41.86 39.60 45.13
Contrast 11.69 13.93 12.75 15.29 23.13
Elastic Transform 20.43 29.64 24.29 27.28 24.40
Pixelate 32.97 41.82 29.86 29.66 46.40
Jpeg Compression 8.00 15.94 9.71 14.79 27.77
Average 17.11 22.72 17.84 20.62 22.38
[…]
keypoint detector : SuperPoint dense
LG LG+Ours GS GS+Ours LoFTR
None (Clean) 67.89 66.98 64.11 61.94 68.88
Gaussian Noise 16.63 28.34 28.57 32.42 24.17
Shot Noise 19.97 32.50 26.23 29.92 26.58
Impulse Noise 23.56 34.74 33.12 36.26 25.67
Defocus Blur 8.29 22.43 12.03 18.17 36.85
Frosted Glass Blur 19.00 32.78 26.39 29.66 40.40
Motion Blur 33.29 39.69 27.88 38.30 39.41
Zoom Blur 20.38 26.23 16.76 25.67 21.60
Snow 39.74 41.07 30.74 33.00 35.91
Frost 47.61 51.78 38.80 41.62 32.97
Fog 62.39 64.08 61.11 60.80 60.74
Brightness 63.40 61.75 62.69 60.04 65.64
Contrast 27.81 33.12 25.86 28.95 30.27
Elastic Transform 47.62 56.66 52.17 55.08 53.61
Pixelate 50.14 56.68 50.33 48.63 63.14
Jpeg Compression 29.92 40.31 40.93 40.61 53.58
Average 33.98 41.48 35.57 38.61 40.70
[…]
Gaussian Noise 8.42 20.10 16.51 20.52 16.68
Shot Noise 10.30 20.52 16.14 19.31 18.14
Impulse Noise 12.88 24.11 21.41 25.35 17.75
Defocus Blur 3.22 17.92 5.95 15.89 24.73
Frosted Glass Blur 8.56 22.42 15.54 18.79 26.87
Motion Blur 18.95 26.22 13.94 24.36 25.92
Zoom Blur 9.81 19.27 8.95 17.68 15.03
Snow 22.94 26.90 17.27 20.56 23.94
Frost 31.58 37.80 24.13 27.95 23.07
Fog 44.49 45.00 42.65 41.24 44.04
Brightness 46.14 45.24 43.91 41.63 48.85
Contrast 14.98 17.15 14.46 17.53 21.09
Elastic Transform 29.34 39.29 32.27 35.11 36.41
Pixelate 31.51 38.02 31.82 30.10 46.26
Jpeg Compression 15.86 27.55 22.24 21.87 37.35
Average 20.60 28.50 21.81 25.19 28.41

### v3

@@ B.2 Robustness against environmental changes
keypoint detector : SuperPoint dense matcher
LG LG+Ours GS GS+Ours LoFTR ELoFTR
None (Clean) 67.89 66.98 64.11 61.94 68.88 72.18
Gaussian Noise 27.44 37.28 28.64 35.11 20.07 20.32
Shot Noise 27.07 39.79 27.81 33.42 24.27 23.89
Impulse Noise 28.70 37.25 29.10 34.44 22.27 23.75
Defocus Blur 18.58 24.51 30.12 33.78 43.27 40.33
Frosted Glass Blur 17.24 27.61 29.16 31.75 39.21 30.43
Motion Blur 27.08 34.85 27.98 35.46 40.68 38.01
Zoom Blur 12.50 20.23 12.64 20.72 12.01 13.66
Snow 18.35 12.56 12.46 10.49 18.94 24.27
Frost 20.86 18.75 17.64 15.48 11.15 15.67
Fog 56.32 62.62 55.14 57.24 53.05 59.85
Brightness 61.10 63.95 58.75 56.51 61.88 67.22
Contrast 24.33 23.10 24.31 27.45 38.57 44.78
Elastic Transform 36.76 47.02 41.53 45.52 41.27 42.79
Pixelate 51.60 60.66 47.60 47.32 62.55 64.85
JPEG Compression 16.60 27.73 22.35 27.64 43.49 43.50
Average 29.64 35.86 31.02 34.16 35.51 36.89
@@
None (Clean) 50.43 49.71 45.92 43.11 52.52 56.38
Gaussian Noise 14.70 24.44 14.63 21.53 9.85 9.57
Shot Noise 13.71 23.44 13.44 18.06 12.21 11.78
Impulse Noise 15.65 21.25 16.90 20.64 10.92 11.77
Defocus Blur 8.22 11.69 16.59 20.64 27.77 25.67
Frosted Glass Blur 7.17 18.09 14.58 16.12 24.59 16.17
Motion Blur 14.72 20.72 15.26 22.34 25.37 27.82
Zoom Blur 5.38 8.82 6.02 14.16 5.17 8.42
Snow 9.48 8.32 5.92 3.87 9.91 13.79
Frost 11.92 11.29 9.92 7.72 6.09 8.75
Fog 39.01 43.43 35.88 37.58 36.94 48.00
Brightness 43.60 47.98 41.86 39.60 45.13 50.73
Contrast 11.69 13.93 12.75 15.29 23.13 29.26
Elastic Transform 20.43 29.64 24.29 27.28 24.40 25.35
Pixelate 32.97 41.82 29.86 29.66 46.40 48.41
JPEG Compression 8.00 15.94 9.71 14.79 27.77 27.45
Average 17.11 22.72 17.84 20.62 22.38 24.20
Table B: The pose accuracy (AUC) at the maximum angular error of $10\tcdegree$ and $5\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [56] under common corruptions. Only one of the input images is corrupted and the other is a clean image. Our method largely improves LightGlue (LG) [27] and GlueStick (GS) [38] for most categories and the average AUC. The results of dense matchers, LoFTR [56], and Efficient LoFTR (ELoFTR) [64] are shown for reference. Our method improves the robustness of sparse matchers with a marginal decrease in clean accuracy.
@@
keypoint detector : SuperPoint dense matcher
LG LG+Ours GS GS+Ours LoFTR ELoFTR
None (Clean) 67.89 66.98 64.11 61.94 68.88 72.18
Gaussian Noise 16.63 28.34 28.57 32.42 24.17 27.01
Shot Noise 19.97 32.50 26.23 29.92 26.58 29.11
Impulse Noise 23.56 34.74 33.12 36.26 25.67 28.67
Defocus Blur 8.29 22.43 12.03 18.17 36.85 31.36
Frosted Glass Blur 19.00 32.78 26.39 29.66 40.40 34.59
Motion Blur 33.29 39.69 27.88 38.30 39.41 37.15
Zoom Blur 20.38 26.23 16.76 25.67 21.60 21.63
Snow 39.74 41.07 30.74 33.00 35.91 43.33
Frost 47.61 51.78 38.80 41.62 32.97 38.00
Fog 62.39 64.08 61.11 60.80 60.74 65.26
Brightness 63.40 61.75 62.69 60.04 65.64 69.24
Contrast 27.81 33.12 25.86 28.95 30.27 40.35
Elastic Transform 47.62 56.66 52.17 55.08 53.61 56.20
Pixelate 50.14 56.68 50.33 48.63 63.14 66.22
JPEG Compression 29.92 40.31 40.93 40.61 53.58 57.23
Average 33.98 41.48 35.57 38.61 40.70 43.02
@@
None (Clean) 50.43 49.71 45.92 43.11 52.52 56.38
Gaussian Noise 8.42 20.10 16.51 20.52 16.68 18.54
Shot Noise 10.30 20.52 16.14 19.31 18.14 19.72
Impulse Noise 12.88 24.11 21.41 25.35 17.75 19.55
Defocus Blur 3.22 17.92 5.95 15.89 24.73 21.55
Frosted Glass Blur 8.56 22.42 15.54 18.79 26.87 22.61
Motion Blur 18.95 26.22 13.94 24.36 25.92 28.77
Zoom Blur 9.81 19.27 8.95 17.68 15.03 17.16
Snow 22.94 26.90 17.27 20.56 23.94 31.62
Frost 31.58 37.80 24.13 27.95 23.07 29.96
Fog 44.49 45.00 42.65 41.24 44.04 51.15
Brightness 46.14 45.24 43.91 41.63 48.85 52.85
Contrast 14.98 17.15 14.46 17.53 21.09 27.79
Elastic Transform 29.34 39.29 32.27 35.11 36.41 38.80
Pixelate 31.51 38.02 31.82 30.10 46.26 49.92
JPEG Compression 15.86 27.55 22.24 21.87 37.35 40.35
Average 20.60 28.50 21.81 25.19 28.41 31.36